OpenStudy (anonymous):

Find all the real cube roots of the number 1.) 1000/27 Find all the real square roots 2.) 0.0625 Find all the real fourth roots 3.) 625 Simplify each radical expressions. Use absolute value symbols when needed. 4.) 3 Radical 64x^9/343

OpenStudy (fibonaccichick666):

hint, break every term into their prime factorizations to start

OpenStudy (anonymous):

How do i do that?

OpenStudy (anonymous):

1. 10/3 2. +0.25 and -0.25 are the real square roots

OpenStudy (anonymous):

3. ± 5

OpenStudy (anonymous):

plus or minus 5 is number 3

OpenStudy (anonymous):

How did you get 1 and 2?

OpenStudy (anonymous):

4. 4x^9/343

OpenStudy (anonymous):

Can you explain how you got the answer to that, please?

OpenStudy (triciaal):

2) 625 * 10^-4 take the square root = 25 * 10^-2 = 0.25

OpenStudy (triciaal):

square root of the square root is the same as the fourth root 4th root of 625 = square root of 25 = 5 therefore the 4th root of 625= 5

OpenStudy (fibonaccichick666):

So, may I do a quick assistance here: @mathhasproblems a prime factorization is where you break a number down until you can't anymore, for example 21=3*7 or 45=9*5=3*3*5

OpenStudy (fibonaccichick666):

for square roots, square implies 2, so you need exactly 2 of the same number in order to pull the number out, for a cube root, cube implies 3, so you need exactly 3 of a number to pull the number out, then it continues

OpenStudy (triciaal):

4) 343 can be written as 7^3 64 = 8^2 and 9 = 3^2

OpenStudy (fibonaccichick666):

here is a quick reference on both topics: http://www.purplemath.com/modules/factnumb.htm http://www.purplemath.com/modules/radicals.htm

OpenStudy (triciaal):

 did you mean the cube root for number 4?